# Verti-Selector: Automatic Curriculum Learning for Wheeled Mobility on Vertically Challenging Terrain

### v1

@@ Abstract
Reinforcement Learning (RL) has the potential to enable extreme off-road mobility by circumventing complex kinodynamic modeling, planning, and control by simulated end-to-end trial-and-error learning experiences. However, most RL methods are sample-inefficient when training in a large amount of manually designed simulation environments and struggle at generalizing to the real world. To address these issues, we introduce Verti-Selector (VS), an automatic curriculum learning framework designed to enhance learning efficiency and generalization by selectively sampling training terrain. VS prioritizes vertically challenging terrain with higher Temporal Difference (TD) errors when revisited, thereby allowing robots to learn at the edge of their evolving capabilities. By dynamically adjusting the sampling focus, VS significantly boosts sample efficiency and generalization within the VW-Chrono simulator built on the Chrono multi-physics engine. Furthermore, we provide simulation and physical results using VS on a Verti-4-Wheeler platform. These results demonstrate that VS can achieve 23.08% improvement in terms of success rate by efficiently sampling during training and robustly generalizing to the real world.
@@ I Introduction
Figure 1: Verti-Selector can selectively sample training terrain based on future learning potential in the VW-Chrono simulator to improve RL sample efficiency and generalization.
Reinforcement learning (RL) offers a promising alternative by enabling robots to learn end-to-end motion policies directly from simulated trial-and-error experiences [10]. By circumventing the need for explicit modeling, planning, and control, RL has the potential to achieve more robust and adaptive off-road navigation. Learning from a high-precision physics model in a simulator with RL in advance can also alleviate onboard computation during deployment.
@@
To push the boundaries of off-road wheeled mobility on vertically challenging terrain, we develop a novel Automatic Curriculum Learning (ACL) method, Verti-Selector (VS), which leverages differences in learning potential across various terrain produced by PCG to enhance both sample efficiency and generalization of RL. VS works in a set of PCG enviornments in a simulator, VW-Chrono, within the Chrono multi-physics simulation engine [15]. Throughout the training process, VS continuously assesses and updates scores that gauge the RL agent’s learning potential on each terrain, taking into account its evolving capability and the Temporal Difference (TD) errors observed from the most recent trajectory sampled from that specific terrain. RL policies efficiently learned with VS in VW-Chrono for navigating vertically challenging terrain can then be deployed onto a physical Verti-4-Wheeler (V4W) platform [16], showing superior real-world generalizability. In summary, the contributions of this work are threefold:
Figure 2: Verti-Selector Overview: The next training terrain $\mathcal{T}_{i}$ is sampled from the training distribution over the PCG-produced terrain set based on priorities determined by evaluation scores of the current policy $\pi$. A trajectory $\tau$ on this terrain $\mathcal{T}_{i}$ is used to update the training distribution.
@@
  • We propose Verti-Selector (VS) (Fig. 2), an ACL framework that samples training terrain based on estimates of future learning potential.
  • Extensive simulation and hardware experiments demonstrate that our approach significantly enhances navigation performance compared against a manually designed curriculum, vanilla RL, and two baseline approaches.
@@ II-A Off-Road Mobility
Considering the limitations of classical approaches, learning off-road mobility has emerged as a promising alternative avenue [22], such as learning end-to-end policies [23], semantic perception [24, 25, 26, 27, 28], kinodynamic models [29, 30, 31, 32, 33, 34, 35], parameter adaptation [36, 37, 38, 39, 40], and cost functions [41, 42, 43, 44, 45, 46, 47, 48]. Learning methods, such as RL, can alleviate engineering effort and allow emergent and adaptive behaviors. However, those methods are often data-hungry, either requiring extensive expert demonstration or labeled data for imitation learning or millions of trial-and-error exploration steps using RL. How to generalize learning results to unseen deployment enviornments is also difficult. To tackle those challenges of learning methods, curriculum learning has the potential to improve sample efficiency and generalization by presenting the robot with a sequence of tasks that gradually increase in difficulty. VS is based on RL guided by an efficient curriculum for wheeled mobility on vertically challenging terrain.
@@ III Method
In this section, we introduce the VW-Chrono simulator (Sec. III-A) and its corresponding PCG environments (Sec. III-B). We also present our RL problem formulation for vertically challenging terrain in Sec. III-C and sample efficient ACL framework, Verti-Selector (VS), in Sec. III-D, which considers robot’s future learning potential on different terrain.
@@ III-A VW-Chrono
To create a realistic simulation environment for vertically challenging terrain, we first collect elevation map data [57] using our physical Verti-4-Wheeler (V4W) on a custom-built indoor testbed. This testbed consists of hundreds of randomly distributed and stacked rocks and boulders, with an average size of 30cm, matching the scale of the V4W. The test course measures 3.1$\times$1.3m, with the highest elevation reaching up to 0.5m, more than twice the height of the vehicle (Fig. 1 middle).
@@ III-B PCG Environments
III-C MDP for Vertically Challenging Terrain
In this work, we formulate the off-road navigation task for wheeled robots as a Markov Decision Process (MDP) characterized by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,r)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ denotes the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathbb{P}(\mathcal{S})$ signifies the transition probability function, $\gamma\in[0,1)$ is the discount factor, and $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function. We employ RL to learn a policy $\pi:\mathcal{S}\to\mathcal{A}$ that maps states $s\in\mathcal{S}$ to actions $a\in\mathcal{A}$, enabling the robot to navigate vertically challenging terrain while avoiding pitfalls such as getting stuck or rolling over, ultimately guiding it to reach the designated goal. The objective is to maximize the expected cumulative discounted reward:
$J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],$ (2)
where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots)$ represents a trajectory sampled from the policy $\pi$.
The state space $\mathcal{S}$ includes the angular difference between the vehicle and goal heading, current vehicle velocity, and a low-dimensional representation of the elevation map patch underneath the robot obtained using a Sliced-Wasserstein Autoencoder (SWAE) [58]. The action space $\mathcal{A}$ consists of the desired linear speed and steering angle, which will be tracked by a low-level PID controller. We employ Proximal Policy Optimization (PPO) [59] as the RL algorithm to learn the policy, considering the continuous action space.
@@ III-C MDP for Vertically Challenging Terrain
By carefully designing the MDP and reward function, we aim to learn a robust policy that enables a wheeled robot to navigate vertically challenging terrain efficiently and safely. The learned policy is then used as a foundation for our ACL approach, which further enhances RL sample efficiency and the robot’s performance and generalization capabilities.
III-D Verti-Selector (VS)
Verti-Selector (VS), depicted in Fig. 2, maintains a dynamic, non-parametric sampling distribution $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ over the set of PCG-generated training terrain $\mathcal{D}_{\text{train}}$, favoring terrain with higher learning potential. Specifically, throughout the training process, VS updates $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ according to a heuristic score that assigns greater weight to terrain with higher estimated learning potential based on the robot’s past experiences. VS maintains two arrays, $\mathbf{u}\in\mathbb{R}^{|\mathcal{D}_{\text{train}}|}$ and $\mathbf{v}\in\mathbb{N}^{|\mathcal{D}_{\text{train}}|}$, where $u_{i}$ stores the score for terrain $\mathcal{T}_{i}$ and $v_{i}$ keeps track of the episode count at which $\mathcal{T}_{i}$ was last sampled. After each episode, VS updates $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ by computing a mixture of two distributions: $q_{u}(\mathcal{T}|\mathcal{D}_{\text{train}})$, based on the terrain scores, and $q_{v}(\mathcal{T}|\mathcal{D}_{\text{train}})$, based on the elapsed time since each terrain was last sampled:
@@ III-D1 Terrain Scoring Mechanism
III-D2 Outdated-Aware Prioritization
To prevent the terrain scores from becoming outdated and to encourage revisiting previously encountered terrain, VS incorporates an outdated-aware prioritization scheme. The outdated-based sampling distribution $q_{v}(\mathcal{T}_{i}|\mathcal{D}_{\text{train}})$ is defined as:
@@ III-D2 Outdated-Aware Prioritization
By combining the score-based and outdated-aware prioritization schemes, VS effectively balances the exploration of terrain with high learning potential and the exploitation of acquired knowledge, leading to more efficient and effective learning in vertically challenging environments.
@@ IV-A Simulation Setup
IV-A1 State Space
The state space consists of the components mentioned in Sec. III-C. The angular difference between the vehicle and goal heading is normalized to $[-1,1]$, and the vehicle velocity is normalized from $[-4,4]$ m/s to $[-1,1]$. The $64\times 64$ elevation map is processed by the SWAE, which reduces it to a $64\times 1$ latent vector. This vector is further compressed to a $16\times 1$ vector using a fully connected neural network. Finally, the compressed elevation map is concatenated with two scalar values, resulting in an $18\times 1$ state vector.
@@ V-A Baselines
VS is compared against four baseline methods: Optimistic Planner (OP), Naive Planner (NP), Vanilla RL (VR), and a Manually-designed Curriculum (MC) [61].
@@ V-B1 Training Performance
The training curve also demonstrates its superior generalization ability by converging to the highest success rate among all methods. By dynamically adjusting the difficulty, VS systematically exposes the robot to a diverse range of terrain features. This diversity helps the robot learn a more comprehensive and flexible policy that can better handle the variability and uncertainty of unseen environments. In contrast, the fixed heuristic of the planners (OP and NP), the lack of a curriculum of VR, and the predefined curriculum of MC may overly specialize the robot to specific terrain types, limiting their generalization to novel test terrain.
@@ VI Conclusions
One of the limitations of VS is its focus on a specific set of terrain geometry and vehicle configurations, which may limit its generalizability to more diverse off-road scenarios. Future work can focus on extending VS to incorporate varioius terrain semantics and vehicle types, as well as integrating with other learning paradigms such as imitation learning.

### v3

# VertiSelector: Automatic Curriculum Learning for Wheeled Mobility on Vertically Challenging Terrain

@@ Abstract
[…]
@@ I Introduction
Figure 1: VertiSelector can selectively sample training terrain based on future learning potential in the VW-Chrono simulator to improve RL sample efficiency and generalization.
Reinforcement learning (RL) offers a promising alternative by enabling robots to learn end-to-end motion policies directly from simulated trial-and-error experiences [10]. By circumventing the need for explicit modeling, planning, and control, RL has the potential to achieve more robust and adaptive off-road navigation. Learning from a high-precision physics model in a simulator with RL in advance can also alleviate simulation-to-reality (sim2real) gap during deployment.
@@
To push the boundaries of off-road wheeled mobility on vertically challenging terrain, we develop a novel Automatic Curriculum Learning (ACL) method, VertiSelector (VS), which leverages differences in learning potential across various terrain produced by PCG to enhance both sample efficiency and generalization of RL. VS works in a set of PCG enviornments in a simulator, VW-Chrono, within the Chrono multi-physics simulation engine [15]. Throughout the training process, VS continuously assesses and updates scores that gauge the RL agent’s learning potential on each terrain, taking into account its evolving capability and the Temporal Difference (TD) errors observed from the most recent trajectory sampled from that specific terrain. RL policies efficiently learned with VS in VW-Chrono for navigating vertically challenging terrain can then be deployed onto a physical Verti-4-Wheeler (V4W) platform [16], showing superior real-world generalizability. In summary, the contributions of this work are threefold:
Figure 2: VertiSelector Overview: The next training terrain $\mathcal{T}_{i}$ is sampled from the training distribution over the PCG-produced terrain set based on priorities determined by evaluation scores of the current policy $\pi$. A trajectory $\tau$ on this terrain $\mathcal{T}_{i}$ is used to update the training distribution.
@@
  • We propose VertiSelector (VS) (Fig. 2), an ACL framework that samples training terrain based on estimates of future learning potential.
  • Extensive simulation and hardware experiments demonstrate that our approach significantly enhances navigation performance compared against a manually designed curriculum, vanilla RL, a hybrid (classical and learning) method and two classical baseline approaches.
@@ II-A Off-Road Mobility
Considering the limitations of classical approaches, learning off-road mobility has emerged as a promising alternative avenue [22], such as learning end-to-end policies [23], semantic perception [24, 25, 26, 27, 28], kinodynamic models [29, 30, 31, 32, 33, 34], parameter adaptation [35, 36, 37, 38, 39], and cost functions [40, 41, 42, 43, 44, 45, 46, 47]. Learning methods, such as RL, can alleviate engineering effort and allow emergent and adaptive behaviors. However, those methods are often data-hungry, either requiring extensive expert demonstration or labeled data for imitation learning or millions of trial-and-error exploration steps using RL. How to generalize learning results to unseen deployment enviornments is also difficult. To tackle those challenges of learning methods, curriculum learning has the potential to improve sample efficiency and generalization by presenting the robot with a sequence of tasks that gradually increase in difficulty. VS is based on RL guided by an efficient curriculum for wheeled mobility on vertically challenging terrain.
@@ III Method
In this section, we introduce the VW-Chrono simulator (Sec. III-A) and its corresponding PCG environments (Sec. III-B). We also present our RL problem formulation for vertically challenging terrain in Sec. III-C and sample efficient ACL framework, VertiSelector (VS), in Sec. III-D, which considers robot’s future learning potential on different terrain.
@@ III-A VW-Chrono
To create a realistic simulation environment for vertically challenging terrain, we first collect elevation map data [56] using our physical Verti-4-Wheeler (V4W) on a custom-built indoor testbed. This testbed consists of hundreds of randomly distributed and stacked rocks and boulders, with an average size of 30cm, matching the scale of the V4W. The test course measures 3.1$\times$1.3m, with the highest elevation reaching up to 0.5m, more than twice the height of the vehicle (Fig. 1 middle). We generate a $150\times 150$ synthetic grayscale elevation map as part of state space to represent real terrain elevation distribution.
@@ III-B PCG Environments
III-C POMDP for Vertically Challenging Terrain
In this work, we formulate the off-road navigation task for wheeled robots as a Partially-Observable Markov Decision Process (POMDP) characterized by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\gamma,r,\mathcal{O},\Omega)$, where $\mathcal{S}$ represents the complete state space, $\mathcal{A}$ denotes the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathbb{P}(\mathcal{S})$ signifies the transition probability function, $\gamma\in[0,1)$ is the discount factor, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $\mathcal{O}$ is the observation space consisting of local terrain and current vehicle information, and $\Omega:\mathcal{S}\times\mathcal{A}\to\mathbb{P}(\mathcal{O})$ is the observation probability function that maps states and actions to observations. We employ RL to learn a policy $\pi:\mathcal{O}\to\mathcal{A}$ that maps observations $o\in\mathcal{O}$ to actions $a\in\mathcal{A}$, enabling the robot to navigate vertically challenging terrain while avoiding pitfalls such as getting stuck or rolling over, ultimately guiding it to reach the designated goal. The objective is to maximize the expected cumulative discounted reward:
$J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(o_{t},a_{t})\right],$ (2)
where $\tau=(o_{0},a_{0},o_{1},a_{1},\ldots)$ represents a trajectory sampled from the policy $\pi$.
To be specific, our observation space $\mathcal{O}$ includes the angular difference between the vehicle and goal heading, current vehicle velocity, and a low-dimensional representation of the elevation map patch underneath the robot obtained using a Sliced-Wasserstein Autoencoder (SWAE) [57]. The action space $\mathcal{A}$ consists of the desired linear speed and steering angle, which will be tracked by a low-level PID controller. We employ Proximal Policy Optimization (PPO) [58] as the RL algorithm to learn the policy, considering the continuous action space.
@@ III-C POMDP for Vertically Challenging Terrain
By carefully designing the POMDP and reward function, we aim to learn a robust policy that enables a wheeled robot to navigate vertically challenging terrain efficiently and safely. The learned policy is then used as a foundation for our ACL approach, which further enhances RL sample efficiency and the robot’s performance and generalization capabilities.
III-D VertiSelector (VS)
VertiSelector (VS), depicted in Fig. 2, maintains a dynamic, non-parametric sampling distribution $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ over the set of PCG-generated training terrain $\mathcal{D}_{\text{train}}$, favoring terrain with higher learning potential. Specifically, throughout the training process, VS updates $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ according to a heuristic score that assigns greater weight to terrain with higher estimated learning potential based on the robot’s past experiences. VS maintains two arrays, $\mathbf{u}\in\mathbb{R}^{|\mathcal{D}_{\text{train}}|}$ and $\mathbf{v}\in\mathbb{N}^{|\mathcal{D}_{\text{train}}|}$, where $u_{i}$ stores the score for terrain $\mathcal{T}_{i}$ and $v_{i}$ keeps track of the episode count at which $\mathcal{T}_{i}$ was last sampled. After each episode, VS updates $q(\mathcal{T}|\mathcal{D}_{\text{train}})$ by computing a mixture of two distributions: $q_{u}(\mathcal{T}|\mathcal{D}_{\text{train}})$, based on the terrain scores, and $q_{v}(\mathcal{T}|\mathcal{D}_{\text{train}})$, based on the elapsed time since each terrain was last sampled:
@@ III-D1 Terrain Scoring Mechanism
III-D2 Staleness-Aware Prioritization
To prevent the terrain scores from becoming outdated and to encourage revisiting previously encountered terrain, VS incorporates a staleness-aware prioritization scheme. The staleness-based sampling distribution $q_{v}(\mathcal{T}_{i}|\mathcal{D}_{\text{train}})$ is defined as:
@@ III-D2 Staleness-Aware Prioritization
By combining the score-based and staleness-aware prioritization schemes, VS effectively balances the exploration of terrain with high learning potential and the exploitation of acquired knowledge, leading to more efficient and effective learning in vertically challenging environments.
@@ IV-A Simulation Setup
IV-A1 Observation Space
The observation space consists of the components mentioned in Sec. III-C. The angular difference between the vehicle and goal heading is normalized to $[-1,1]$, and the vehicle velocity is normalized from $[-4,4]$ m/s to $[-1,1]$. The $64\times 64$ elevation map is processed by the SWAE, which reduces it to a $64\times 1$ latent vector. This vector is further compressed to a $16\times 1$ vector using a fully connected neural network. Finally, the compressed elevation map is concatenated with two scalar values, resulting in an $18\times 1$ state vector.
@@ V-A Baselines
VS is compared against four baseline methods: Optimistic Planner (OP), Naive Planner (NP), Vanilla RL (VR), Manually-designed Curriculum (MC) [60] and WMVCT [61].
@@
WMVCT is a hybrid method based on a sampling-based motion planner and a decomposed 6-DoF vehicle-terrain dynamics model (bicycle model for x, y, and yaw, elevation map for z, and neural network prediction for roll and pitch). Despite its high efficiency, the decomposition also introduces inaccuracies compared against a full 6-DoF model.
@@ V-B1 Training Performance
The training curve also demonstrates its superior generalization ability by converging to the highest success rate among all methods. By dynamically adjusting the difficulty, VS systematically exposes the robot to a diverse range of terrain features. This diversity helps the robot learn a more comprehensive and flexible policy that can better handle the variability and uncertainty of unseen environments. In contrast, the fixed heuristic of the planners (OP and NP), the decomposed 6-DoF vehicle-terrain dynamics model of WMVCT, the lack of a curriculum of VR, and the predefined curriculum of MC may overly specialize the robot to specific terrain types, limiting their generalization to novel test terrain.
@@ VI Conclusions and Limitations
One of the limitations of VS is its focus on a specific set of terrain geometry and vehicle configurations, which may limit its generalizability to more diverse off-road scenarios. Despite the effectiveness of VW-Chrono, sim2real gap still remains: the framework trains and evaluates using a simulated full-scale vehicle, while real-world validation employs a scaled RC car. This discrepancy in vehicle scale and dynamics may undermine the transferability of results to practical scenarios. And it is also important to note that real rocky terrain exhibits spatially varying friction and restitution properties. Future work can focus on extending VS to incorporate various terrain semantics and vehicle types, as well as integrating with other learning paradigms such as imitation learning.